Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Monday, November 23, 1998

The following is a brief summary of the environmental monitoring activities that were completed on November 23, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

11/23/98 (Monday)

Segment 1 report for 11/23/98: On Peace Valley Rd., casing for the bore (not yet begun) was being welded at EMP 25.43. Cleaner pigs were run through the segment where hydrostatic testing will take place from EMP's 25.43 to 21.75. Welding took place on Lebec Rd. at the Warbler skip from EMP 21.47 to 21.50. On Tejon Bench trench was excavated to EMP 18.10 and a section of pipe was tied-in. Backfill and compaction took place from EMP 17.14 to 17.19. On Grapevine Rd., the coating crew worked to EMP 15.95, welding and x-ray to 15.90. At Grapevine Pump Station electrical work and pipe fitting continued. Placement of conduit and backfill occurred on both side of the I-5/Grapevine bore (EMP 14.62 to 14.69) and a pull box was installed at EMP 14.63. The bore operations continued at the California Aqueduct (12.28), the stream crossing #4 bore (9.03) and the Hwy. 166 bore (1.18). Cable was pulled from EMP 8.62 to 6.57. On Hwy. 166, welding and x-ray took place at EMP .56. Flooring was being installed in the office at Emidio Pump Station and PG&E was installing power lines to the pump station.

Segment 2B (Castaic to Sylmar): Report for 11-23-98. Work continued on the electrical connections for the vaults at Romero Cyn and Rye Cyn Rd. Drilling was completed at the Santa Clara river and crews began a general clean up of the drilling site. In addition, crews anticipate pulling the 1300' casing on Tuesday. Further, pipe was removed from Feedmill Rd and along The Old Rd at Valencia Blvd. South of Weldon Cyn Rd, trenching crews completed 350' of ditch while bending and stringing crews worked nearby. Welding, x-ray and coating crews progressed directly behind the pipe crews at EMP 72.44. In addition, several hundred feet of slurry was poured at locations 300' north of Sierra Hwy while asphalt was completed up to Sierra Hwy.

Inspection Summary

In the past 24 hours, 2 inspection event(s) were reported for 2 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page